# Vector Cost Bimatrix Games with Applications to Autonomous Racing

### v1

@@ Abstract
We formulate a vector cost alternative to the scalarization method for weighting and combining multi-objective costs. The algorithm produces solutions to bimatrix games that are simultaneously pure, unique Nash equilibria and Pareto optimal with guarantees for avoiding worst case outcomes. We achieve this by enforcing exact potential game constraints to guide cost adjustments towards equilibrium, while minimizing the deviation from the original cost structure. The magnitude of this adjustment serves as a metric for differentiating between Pareto optimal solutions. We implement this approach in a racing competition between agents with heterogeneous cost structures, resulting in fewer collision incidents with a minimal decrease in performance. Code is available at https://github.com/toazbenj/race\_simulation.
@@
thanks: This work was supported by the National Science Foundation CAREER Award ECCS-2236537.
@@ 3 Proposed Approach and Main Results
This leads us to the following optimization problem. The input is the desired location of the minimum of $\phi$ given by entry $(r,c)$ and the cost matrices $A_{1}$ and $C_{2}$.
@@
In this optimization problem, the first two constraints model the requirement that the matrix $\phi$ is a potential function for the bimatrix game $A_{1}+E,C_{2}$. The final constraint models the requirement that the global minimum of $\phi$ is exactly at the specified choice location $(r,c)$.
@@
Now our approach, summarized in Algorithm 1, computes the optimal error matrices for every choice of candidate global minima $(r,c)$ in $\phi$. Note that should the algorithm fail to find a suitable $E$ using any $(r,c)$, player 1 will default back to the scalarization method.
@@
For a given value of the costs $C_{2}$ for player 2, Equation 8 produces finite $E$ such that bimatrix game $\tilde{A}_{1},C_{2}$ produces potential $\phi$ with a global minimum at chosen position $(r,c)$ if fixed costs $C_{2}$ adhere to the following conditions. For entry $d_{rj}\in D^{row}(C_{2})$:
@@
Notice that for all rows other than row $r$, we can select a $k_{i}$ such that all of the row entries are greater than zero. For row $r$ however, entries $d_{rj}$ must be already be in a range that results in positive values across every entry that is not the minimum. If $d_{rj}>0$ for entries to the left of column $c$, the value will be negative and invalidate this constraint. Similarly, this will be the case if $d_{rj}<0$ for values to the right of column $c$. Therefore, in order to select global minimum $(r,c)$, the row pairwise differences of $C_{2}$ must adhere to the constraints described in Theorem 2. ∎
@@ 4 Numerical Comparison with Baseline
Unfortunately in this case we cannot pick a pair $(\theta_{1},\theta_{2})$ that results in an equilibria at $(2,2)$. Reducing the weight of $\theta_{1}$ to intermediate values does not change the security policy outcome as long as $\theta_{1}>\theta_{2}$. Security policies will always be Nash equilibria within a single game type, but result in worst-case outcomes in the other game.
@@ 5 Application to autonomous racing
Scenario Passes Collisions Attacker Off Track
@@
Scenario Attacker Lead Time (%) Attacker Avg. Laps Defender Avg. Laps

### v3

@@ Abstract
We formulate a vector cost alternative to the scalarization method for weighting and combining multi-objective costs. The algorithm produces solutions to bimatrix games that are simultaneously pure, unique Nash equilibria and Pareto optimal with guarantees for avoiding worst case outcomes. We achieve this by enforcing exact potential game constraints to guide cost adjustments towards equilibrium, while minimizing the deviation from the original cost structure. The magnitude of this adjustment serves as a metric for differentiating between Pareto optimal solutions. We implement this approach in a racing competition between agents with heterogeneous cost structures, resulting in fewer collision incidents with a minimal decrease in performance. Code is available at https://github.com/toazbenj/race_simulation.
@@
thanks: This work was supported by the National Science Foundation CAREER Award ECCS-2236537. Special thanks also to the Spartan Autonomous Racing Club at Michigan State.
@@ 3 Proposed Approach and Main Results
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 4 Numerical Comparison with Baseline
Unfortunately in this case we cannot pick a pair $(\theta_{1},\theta_{2})$ that results in an equilibria at (2,2)(2,2)( 2 , 2 ). Reducing the weight of $\theta_{1}$ to intermediate values does not change the security policy outcome as long as $\theta_{1}>\theta_{2}$. Security policies will always be Nash equilibria within a single game type, but result in worst-case outcomes in the other game.
@@ 5 Application to autonomous racing
Scenario
Passes
Collisions
Attacker
Off Track
@@
Scenario
Attacker Lead
Time (%)
Attacker
Avg. Laps
Defender
Avg. Laps